OpenStudy (anonymous):

In a toy factory 200 wooden solid cylinders 7cm long and 35mm in diameter have to be painted. What is the total surface area in cm sq?

OpenStudy (anonymous):

Are you looking for the surface area of all 200 combined? Or just one cylinder?

OpenStudy (anonymous):

Combined

OpenStudy (anonymous):

Hint: The surface area of a cylinder is: \[2 \pi*r ^{2}+2 \pi*r*h\] find the surface area of single cylinder and multiply by 200

OpenStudy (anonymous):

Okay, well the formula to find the surface area is (2*pi*r^2)+(2*pi*r*h) r= radius h= height so you would multiply 2 times 3.14(pi) times the radius and then square it. Now the question only gives you the diameter whereas the radius is half of the diameter so you would divide 35 by 2 and you would get 17.5. So thats the first part of your equation. Then the next part would be 2 times 3.14(pi) times the radius (17.5) times the height which is 7 and you need to convert the 7 cm down to mm which would be 70mm. So your overall equation would look like: (2*pi*17.5^2)+(2*pi*17.5*70). Now you can solve this on your own and tell me the answer that you get.

OpenStudy (anonymous):

Then multiply by 200

hero (hero):

@n_deatz, @stgreen posted a more appropriate response. I highly recommend it for this particular user.

OpenStudy (anonymous):

@Hero thank you for your opinion. I was only trying to help since this is the way my teachers taught me. But thank you. :)

hero (hero):

Yes, but there's a fundamental thing about teaching....it's called guidance. The only thing I see is that you posted an answer. I don't see any guidance being administered here.

hero (hero):

If you really wanted to guide her, you would have asked her what the specific question was she had concerning this problem to figure out exactly what it is that is confusing her. You could even ask what she has done so far to solve the problem. You could have shared the formula with her if she didn't know it and discussed what the variables meant and what values they represent and what to solve for.

hero (hero):

@n_deatz

OpenStudy (anonymous):

I understand. Thank you. @Hero